OpenStudy (anonymous):

A rancher wants to fence in an area of 1000000 square feet in a rectangular field and then divide it in half with a fence down the middle parallel to one side. What is the shortest length of fence that the rancher can use

OpenStudy (anonymous):



OpenStudy (anonymous):

you have \(LW=1,000,000\) so \(W=\frac{1,000,000}{L}\) and you want to minimize \[3L+W=3L+\frac{1,000,000}{L}\]

OpenStudy (anonymous):

this gives you \[f(L)=3L+\frac{1,000,000}{L}\] to find the minimum, take the derivative, find the critical points by setting the derivative equal to zero, and solve for \(L\)

OpenStudy (anonymous):

you good from there?